# Notions of enriched purity

### v2

@@ 6. Purity over quantale-valued metric spaces
$x+\bigwedge\limits_{j\in J}y_{j}=\bigwedge\limits_{j\in J}x+y_{j}$
for every $x,y_{j}\in Q$, $j\in J$. A totally above relation $a\succ b$ means that for every subset $S\subseteq Q$ such that $a\geq\bigwedge S$ implies that there exists $s\in S$ such that $b\geq s$. Let $\Uparrow(a)=\{x\in Q|x\succ a\}$. A value quantale is a quantale $Q$ such that $a=\bigwedge\Uparrow(a)$ for every $a\in Q$ and, moreover, $a\wedge b\succ 0$ whenever $a,b\succ 0$ (see [8, 6]).
@@
A map $f:X\to Y$ is nonexpanding if $d(fx,fy)\leq d(x,y)$ for all $x,y\in X$. The resulting category is denoted by $Q$-$\operatorname{\bf Met}$.
@@
(2) Ultrametric spaces are $Q$-quantales for $Q=(\{0,\infty],\leq,\vee)$.
@@
(see [8], or [5, 1.2(4)]).
The tensor product $X\otimes Y$ of $Q$-valued metric spaces is $X\times Y$ with the distance
@@
The one-element $Q$-metric space is the tensor unit $I$. The hom-set $[X,Y]$ is a $Q$-metric space with the distance
@@
Remark 6.3.
If $Q=([0,\infty],\leq,+)$ and we drop the symmetry condition (2) in Definition 6.1, we obtain the category of (skeletal) Lawvere metric spaces [19]; that is, the category $[0,\infty]$-$\mathbf{Cat}_{sk}$ of small and skeletal $[0,\infty]$-enriched categories. Moreover, the symmetric monoidal structure on such metric spaces coincides with the one induced from the standard tensor product of enriched categories. The same properties hold for a general value quantale $Q$: every $Q$-metric space is a skeletal $Q$-enriched category (which in addition satisfies the symmetry condition). Moreover, the tensor product described above is the same as the tensor product of $Q$-enriched categories. Note that the monoidal structure on $Q$ induced by addition is symmetric monoidal closed with internal hom [,][-,-][ - , - ] given by
$[x,y]=(y-x):=\bigvee\limits_{z+y\leq x}z$
as for $[0,\infty]$.
@@
Consider a nonexpanding map $f:X\otimes Y\to Z$. Since
@@
the maps $f(x,-):Y\to Y$ are nonexpanding. Since
@@
the map $\tilde{f}:X\to[Y,Z]$ such that $\tilde{f}x=f(x,-)$ is nonexpanding.
@@
the map $\tilde{g}$ is nonexpanding. ∎
@@
A map $f:X\to Y$ is an isometry if $d(fx,fy)=d(x,y)$ for all $x,y\in X$. Following 6.6, isometries are closed under directed colimits in ($Q$-$\operatorname{\bf Met})^{\to}$.
@@
A key property of value quantales is that for every $q\succ 0$ there is $p\succ 0$ such that $q\succ 2p$ (see [8, 2.9]). This makes possible to define a topology on a $Q$-metric space $(X,d)$ as follows (see [8, 4.1]). For $x\in X$ and $q\succ 0$, the open ball $B_{q}(x)$ is the set $\{y\in X|q\succ d(x,y)\}$. Now, a set $U\subseteq X$ is open if for every $x\in U$ there is $q\succ 0$ such that $B_{q}(x)\subseteq U$.
@@
is a square such that $fu\sim_{q}vg$. In this setting the notion of $q$-pushout that we introduce below is relevant: this will play a key role in the proof of Proposition 6.10. Like in [3], $q$-pushouts can be seen as weighted (finite) colimits.
@@
$(1)\Rightarrow(2)$ is straightforward since every $u\in\mathcal{P}(g,f)$ can be completed to a $q$-commutative square as in the definition of weak $(\lambda,\mathcal{E})$-purity, while $(2)\Rightarrow(3)$ is given by Lemma 4.4. $(3)\Rightarrow(1)$. Let $f$ be barely $(\lambda,\mathcal{E})$-pure and consider a $q$-commutative square
@@ 7. Purity over $\omega$-complete posets
with the same notations we used in Example 5.2(4) to define the dense maps as an injectivity class. Then, to give $g\colon A\to B$ and $f\colon A\to C$ in $\mathcal{K}$ is the same as giving an enriched functor $H\colon\mathcal{C}\to\mathcal{K}$, and it is easy to see that the $\omega$-pushout of $g$ along $f$ coincides with the weighted colimit $M*H$.
We are now ready to prove the two notions of purity coincide (like in 6.10):
@@
$(1)\Rightarrow(2)$ follows from Lemma 4.4, and $(2)\Rightarrow(3)$ is trivial since it is enough to apply the bare $(\lambda,\mathcal{E})$-purity of $f$ to the constant family $(u_{i},v_{i})=(u,v)$. $(3)\Rightarrow(1)$. Following the definition of $(\lambda,\mathcal{E})$-purity, we need to consider $g:A\to B$ between $\lambda$-presentable objects, a map $u:A\to K$, and a family $v_{i}\colon B\to L$ such that $v_{i}g$ is a chain and $fu=\bigvee v_{i}g$. Consider now the $\omega$-pushout of $g$ along $1_{A}$ in $\mathcal{K}$; this is an object $D$ together with maps $\bar{g}\colon A\to D$ and $h_{i}\colon B\to D$, for $i\in\omega$, such that $(h_{i}g)_{i\in\omega}$ is a chain and $\bar{g}=\bigvee h_{i}g$. By the universal property of the $\omega$-pushout applied to the tuple $(L,fu,(v_{i})_{i\in\omega})$ there exists a map $\bar{v}\colon D\to L$ making the square
@@ 8. Purity over quasivarieties
In particular, such a $\mathcal{V}$ is also locally $\lambda$-presentable as a closed category in the sense of Kelly [13] (see [16, Remark 4.15]). Following [4, Proposition 30], a regular generator in (1) can be weakened to strong generator.
@@
    $\mathbf{Set}$, where the induced factorization system onduces the standard notion of purity;
@@
In $\mathbf{DGAb}$, consider the chain complex $P_{n}$ which has $\mathbb{Z}$ in degree $n$ and $n-1$, with differential $d_{n}=\textnormal{id}$, and which is trivial in every other degree. In [21] these $P_{n}$ are called $S^{n}L\mathbb{Z}$. To give a map $P_{n}\to A$ in $\mathbf{DGAb}$ is the same as giving an element $x\in A_{n}$ in degree $n$. Then, according to the nomenclature in [21], a $P_{n}$-morphism in a DG-category $\mathcal{K}$ is simply a protomorphism of degree $n$. It is easy to see that the set of all $P_{n}$’s is a regular generator of $\mathbf{DGAb}$ made of finitely presentable and regular projective objects.
@@
The morphisms above were considered in [30] to express every $\mathcal{V}$-category as a category enriched over $[\mathcal{V}^{\textnormal{op}},\mathbf{SET}]$, for an opportune universe enlargement $\mathbf{SET}$ of $\mathbf{Set}$. In particular, the definition below can be as part of the composition rule in $[\mathcal{V}^{\textnormal{op}},\mathbf{SET}]$-categories.
@@
Thanks to this we can talk abut commutative squares involving $P$-morphisms:
@@
thus $\mathcal{V}_{0}(P,\mathcal{Q}(g,f))$ is the standard image factorization of the function $\mathcal{V}_{0}(P,\pi_{1})$. It follows that $\mathcal{V}_{0}(P,\mathcal{Q}(g,f))$ can be described as the set of those $P$-morphisms $(u,P)\colon A\to K$ which can be completed to a commutative square as depicted in the statement. In conclusion, the map $\mathcal{V}_{0}(P,q)$ is surjective if and only if for any $(u,P)$ as above there exists $(t,P)\colon B\to K$ such that $(tg,P)=(u,P)$. ∎
@@
Since maps $P\cdot A\to B$ in $\mathcal{K}$ correspond to morphisms $P\to\mathcal{K}(A,K)$ in $\mathcal{V}$, to give a square as in $(3)$ is the same as giving a commutative square involving $P$-morphisms as in Proposition 8.9 above. Therefore $(1)\Leftrightarrow(3)$ follows immediately. For $(1)\Leftrightarrow(2)$ it is again enough to notice that, acting by transposition, a square as in Proposition 8.9 is the same as a commutative square
@@
If $f\colon K\to L$ is ordinarily $\lambda$-pure then it satisfies point $(3)$ of Proposition 8.10 since copowers by finitely presentable objects of $\mathcal{V}$ are still finitely presentable. Therefore $f$ is $(\lambda,\mathcal{E})$-pure. If the unit $I$ is regular projective, then we can assume it to be an element of $\mathcal{P}$ (we can always take $\mathcal{P}$ to be the set of all the finitely presentable and regular projective objects). Thus, $f\cong I\pitchfork f$ is ordinarily $\lambda$-pure by Proposition 8.10. ∎
@@
Recall from the notion of $\mathcal{E}$-injectivity Section 3. In this context it translates into the following: given a $\mathcal{V}$-category $\mathcal{K}$, an object $X$ is $\mathcal{E}$-injective with respect to $h\colon A\to B$ in $\mathcal{K}$ if
@@
Note that Assumption 5.1 is satisfied, and that a morphism is $(\lambda,\mathcal{E})$-pure if and only if it is barely $(\lambda,\mathcal{E})$-pure Lemma 4.4 since regular epimorphisms are stable under pullbacks in $\mathcal{V}$. Moreover, a map $e$ in $\mathcal{V}$ is a regular epimorphism if and only if $\mathcal{V}_{0}(P,e)$ is surjective for any $P\in\mathcal{P}$ (since $\mathcal{P}$ forms a regular generator), and powers by by $\mathcal{P}$ are $\mathcal{E}$-stable by [16, Remark 4.15]. Thus we can apply Theorem 5.5 for $\mathcal{G}:=\mathcal{P}$ and obtain the desired characterization of $(\lambda,\mathcal{E})$-injectivity classes. ∎
@@
Since $\mathcal{V}$ is a $\lambda$-quasivariety and the unit $I$ is $\lambda$-presentable, we can write it as a $\lambda$-small colimit of $\lambda$-presentable regular projective objects. Thus there exists a $\lambda$-presentable and regular projective object $I^{*}$ together with a regular epimorphism $e\colon I^{*}\twoheadrightarrow I$. We shall now fix such $I^{*}$ and $e$ and, without loss of generality, assume that $I^{*}\in\mathcal{P}$.
@@
Given $I^{*}$ as above, by pre-composing with $e$, every morphism in a $\mathcal{V}$-category $\mathcal{K}$ can be seen as a $I^{*}$-morphism. As it happens in the case of ordinary morphisms (Definition 8.6), $I^{*}$-morphisms can be composed with $P$-morphisms for any $P\in\mathcal{P}$. Indeed, given $P\in\mathcal{P}$, tensoring with $e$ gives a regular epimorphism $e_{P}\colon I^{*}\otimes P\to P$ which, since $P$ is regular projective, splits giving $s_{e}\colon P\to I^{*}\otimes P$ such that $e_{P}s_{P}=1_{P}$. Now given a $P$-morphism $(f,P)\colon X\to Y$ and a $I^{*}$-morphism $(g,I^{*})\colon Y\to Z$, we define their composite as the $P$-morphism
$(gf,P)\colon P\stackrel{{\scriptstyle s_{e}}}{{\longrightarrow}}I^{*}\otimes P\stackrel{{\scriptstyle g\otimes f}}{{\xrightarrow{\hskip 22.76228pt}}}\mathcal{K}(Y,Z)\otimes\mathcal{K}(X,Y)\stackrel{{\scriptstyle\circ}}{{\longrightarrow}}\mathcal{K}(X,Z).$
Finally, note that the composite of a $P$-morphism $(f,P)\colon X\to Y$ with a morphism $g\colon Y\to Z$ is the same as the composite of $(f,P)$ with the $I^{*}$-morphism induced by $g$ through $e$. This is because $s_{P}$ is a section of $e_{P}$.
@@
A morphism $s\colon X\to Y$ in a $\mathcal{V}$-category $\mathcal{K}$ is an $\mathcal{E}$-split monomorphism if and only if there exists $(t,I^{*})\colon Y\to X$ for which $(ts,I^{*})=(1_{X},I^{*})$.
@@
is surjective. By taking an element that is mapped to $(1_{X},I^{*})$ we find $(t,I^{*})\colon Y\to X$ as in the statement. Conversely, suppose that there exist $(t,I^{*})\colon Y\to X$ for which $(ts,I^{*})=(1_{X},I^{*})$. To prove that $s$ is an $\mathcal{E}$-split monomorphism it is enough to show that $\mathcal{V}_{0}(P,\mathcal{K}(s,X))$ is surjective for any $P\in\mathcal{P}$. Given $(f,P)\colon X\to X$ in the codomain, it can be seen as the image through $\mathcal{V}_{0}(P,\mathcal{K}(s,X))$ of $(ft,P)\colon Y\to X$. Thus $\mathcal{K}(s,X)$ is a regular epimorphism and $s$ an $\mathcal{E}$-split monomorphism. ∎
@@
(1) follows from Corollary 5.4 and (2) is a consequence of Proposition 4.9 since monomorphisms are closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$. (3) We apply the same proof of [2, Proposition 2.30] with some changes where the notion of $\mathcal{E}$-split morphism is needed. Given an $\mathcal{E}$-pure morphism $f\colon A\to B$, we can write it as a $\lambda$-filtered colimit in $\mathcal{K}^{\to}$ of maps $f_{i}\colon A_{i}\to B_{i}$ between $\lambda$-presentable objects with connecting morphisms $(u_{i},v_{i})\colon f_{i}\to f$. As usual, all the morphisms $u_{i}$ and $v_{i}$ can be seen as $I^{*}$-morphisms, giving commutative squares as below.
@@
It is shown in [2, Proposition 2.30] that the colimit of the $\bar{f}_{i}$ is still $f$; thus we only need to show that each $\bar{f}_{i}$ is $\mathcal{E}$-split. We can see the identity on $A$ as an $I^{*}$-morphism and hence, acting by transposition, as a map $\bar{1}_{A}\colon A\to I^{*}\pitchfork A$; similarly each $(t_{i},I^{*})$ corresponds to a morphism $\bar{t}_{i}\colon B_{i}\to I^{*}\pitchfork A$. The fact that $(t_{i}f,I^{*})=(u_{i},I^{*})$ means that the square below commutes.

### v3

@@ 6. Purity over quantale-valued metric spaces
$x+\bigwedge\limits_{j\in J}y_{j}=\bigwedge\limits_{j\in J}(x+y_{j})$
for every $x,y_{j}\in Q$, $j\in J$. Given $a,b\in Q$ we say that $a$ is totally above $b$, and write $a\succ b$, if: for every subset $S\subseteq Q$ such that $a\geq\bigwedge S$ there exists $s\in S$ such that $b\geq s$. Let now $\Uparrow(a)=\{x\in Q|x\succ a\}$; a value quantale is a quantale $Q$ such that
$a=\bigwedge\Uparrow(a)$
for every $a\in Q$ and, moreover, $a\wedge b\succ 0$ whenever $a,b\succ 0$ (see [9, 7]).
@@
A map $f:X\to Y$ is non-expanding if $d(fx,fy)\leq d(x,y)$ for all $x,y\in X$. The resulting category is denoted by $Q$-$\operatorname{\bf Met}$.
@@
(2) Ultrametric spaces are $Q$-quantales for $Q=((0,\infty],\leq,\vee)$.
@@
(see [9], or [6, 1.2(4)]).
The tensor product $X\otimes Y$ of $Q$-valued metric spaces is given by the set $X\times Y$ with the distance
@@
The tensor unit $I$ is the one-element $Q$-metric space. The internal hom $[X,Y]$ is the $Q$-metric space on the set $Q$-$\operatorname{\bf Met}(X,Y)$ with distance
@@
Clearly $\otimes$ is symmetric and $I$ is a unit for it. Consider a nonexpanding map $f:X\otimes Y\to Z$. Since
@@
the maps $f(x,-)\colon Y\to Y$ are nonexpanding. Moreover, since
@@
the map $\tilde{f}\colon X\to[Y,Z]$ such that $\tilde{f}x=f(x,-)$ is nonexpanding.
@@
the map $\tilde{g}$ is nonexpanding. This defines a natural bijection between maps $X\otimes Y\to Z$ and $X\to[Y,Z]$. ∎
@@
A map $f:X\to Y$ is an isometry if $d(fx,fy)=d(x,y)$ for all $x,y\in X$. Following 6.5, isometries are closed under directed colimits in ($Q$-$\operatorname{\bf Met})^{\to}$.
@@
A key property of value quantales is that for every $q\succ 0$ there is $p\succ 0$ such that $q\succ 2p$ (see [9, 2.9]). This makes possible to define a topology on a $Q$-metric space $(X,d)$ as follows (see [9, 4.1]). For $x\in X$ and $q\succ 0$, the open ball $B_{q}(x)$ is the set $\{y\in X|q\succ d(x,y)\}$. Now, a set $U\subseteq X$ is open if for every $x\in U$ there is $q\succ 0$ such that $B_{q}(x)\subseteq U$.
@@
is a square such that $fu\sim_{q}vg$. In this setting the notion of $q$-pushout that we introduce below is relevant: this will play a key role in the proof of Proposition 6.9. Like in [3], $q$-pushouts can be seen as weighted ($\aleph_{1}$-small) colimits.
@@
$(1)\Rightarrow(2)$ is straightforward since every $u\in\mathcal{P}(g,f)$ can be completed to a $q$-commutative square as in the definition of weak $(\lambda,\mathcal{E})$-purity, while $(2)\Rightarrow(3)$ is given by Lemma 4.4.
$(3)\Rightarrow(1)$. Let $f$ be barely $(\lambda,\mathcal{E})$-pure and consider a $q$-commutative square
@@ 7. Purity over $\omega$-complete posets
with the same notations we used in Example 5.2(4) to define the dense maps as an injectivity class.
Then, to give $g\colon A\to B$ and $f\colon A\to C$ in $\mathcal{K}$ is the same as giving an enriched functor $H\colon\mathcal{C}\to\mathcal{K}$, and it is easy to see that the $\omega$-pushout of $g$ along $f$ coincides with the weighted colimit $M*H$.
We are now ready to prove the two notions of purity coincide (like in 6.9):
@@
$(1)\Rightarrow(2)$ follows from Lemma 4.4, and $(2)\Rightarrow(3)$ is trivial since it is enough to apply the bare $(\lambda,\mathcal{E})$-purity of $f$ to the constant family $(u_{i},v_{i})=(u,v)$.
$(3)\Rightarrow(1)$. Following the definition of $(\lambda,\mathcal{E})$-purity, we need to consider $g:A\to B$ between $\lambda$-presentable objects, a map $u:A\to K$, and a family $v_{i}\colon B\to L$ such that $v_{i}g$ is a chain and $fu=\bigvee v_{i}g$.
Consider now the $\omega$-pushout of $g$ along $1_{A}$ in $\mathcal{K}$; this is an object $D$ together with maps $\bar{g}\colon A\to D$ and $h_{i}\colon B\to D$, for $i\in\omega$, such that $(h_{i}g)_{i\in\omega}$ is a chain and $\bar{g}=\bigvee h_{i}g$. By the universal property of the $\omega$-pushout applied to the tuple $(L,fu,(v_{i})_{i\in\omega})$ there exists a map $\bar{v}\colon D\to L$ making the square
@@ 8. Purity over quasivarieties
In particular, such a $\mathcal{V}$ is also locally $\lambda$-presentable as a closed category in the sense of Kelly [13] (see [16, Remark 4.15]). Following [4, Proposition 30], a regular generator as in (1) can be weakened to be just a strong generator.
@@
    $\mathbf{Set}$, where the induced factorization system induces the standard notion of purity;
@@
In the monoidal category $\mathbf{DGAb}$ of chain complexes, consider the chain complex $P_{n}$ which has $\mathbb{Z}$ in degree $n$ and $n-1$, with differential $d_{n}=\textnormal{id}$, and which is trivial in every other degree. In [20] these $P_{n}$ are called $S^{n}L\mathbb{Z}$. It is easy to see that to give a morphism $P_{n}\to A$ in $\mathbf{DGAb}$ is the same as specifying an element $x\in A_{n}$ in degree $n$.
Then, according to the nomenclature in [20], a $P_{n}$-morphism in a DG-category $\mathcal{K}$ is a protomorphism of degree $n$. It is easy to see that the set of all $P_{n}$’s is a regular generator of $\mathbf{DGAb}$ made of finitely presentable and regular projective objects (in fact, this is a dense generator by [20, 3.10]).
@@
The morphisms above were considered in [30] to express every $\mathcal{V}$-category as a category enriched over $[\mathcal{V}^{\textnormal{op}},\mathbf{SET}]$, for an opportune universe enlargement $\mathbf{SET}$ of $\mathbf{Set}$. In particular, the definition below can be seen as part of the composition rule in $[\mathcal{V}^{\textnormal{op}},\mathbf{SET}]$-categories.
@@
Thanks to this we can talk about commutative squares involving $P$-morphisms:
@@
thus $\mathcal{V}_{0}(P,\mathcal{Q}(g,f))$ is the standard image factorization of the function $\mathcal{V}_{0}(P,\pi_{1})$. It follows that $\mathcal{V}_{0}(P,\mathcal{Q}(g,f))$ can be described as the set of those $P$-morphisms $(u,P)\colon A\to K$ which can be completed to a commutative square as depicted in the statement.
In conclusion, the map $\mathcal{V}_{0}(P,q)$ is surjective if and only if for any $(u,P)$ as above there exists $(t,P)\colon B\to K$ such that $(tg,P)=(u,P)$. ∎
@@
Since maps $P\cdot A\to B$ in $\mathcal{K}$ correspond to morphisms $P\to\mathcal{K}(A,K)$ in $\mathcal{V}$, to give a square as in $(3)$ is the same as giving a commutative square involving $P$-morphisms as in Proposition 8.9 above. Therefore $(1)\Leftrightarrow(3)$ follows immediately.
For $(1)\Leftrightarrow(2)$ it is again enough to notice that, acting by transposition, a square as in Proposition 8.9 is the same as a commutative square
@@
If $f\colon K\to L$ is ordinarily $\lambda$-pure then it satisfies point $(3)$ of Proposition 8.10 since copowers by finitely presentable objects of $\mathcal{V}$ are still finitely presentable. Therefore $f$ is $(\lambda,\mathcal{E})$-pure.
If the unit $I$ is regular projective, then we can assume it to be an element of $\mathcal{P}$ (we can always take $\mathcal{P}$ to be the set of all the finitely presentable and regular projective objects). Thus, $f\cong I\pitchfork f$ is ordinarily $\lambda$-pure by Proposition 8.10. ∎
@@
Recall the notion of $\mathcal{E}$-injectivity from Section 3. In this context it translates into the following: given a $\mathcal{V}$-category $\mathcal{K}$, an object $X$ is $\mathcal{E}$-injective with respect to $h\colon A\to B$ in $\mathcal{K}$ if
@@
Note that Assumption 5.1 is satisfied, and that a morphism is $(\lambda,\mathcal{E})$-pure if and only if it is barely $(\lambda,\mathcal{E})$-pure Lemma 4.4 since regular epimorphisms are stable under pullbacks in $\mathcal{V}$. Moreover, a map $e$ in $\mathcal{V}$ is a regular epimorphism if and only if $\mathcal{V}_{0}(P,e)$ is surjective for any $P\in\mathcal{P}$ (since $\mathcal{P}$ forms a regular generator), and powers by $\mathcal{P}$ are $\mathcal{E}$-stable by [16, Remark 4.15].
Thus we can apply Theorem 5.5 for $\mathcal{G}:=\mathcal{P}$ and obtain the desired characterization of $(\lambda,\mathcal{E})$-injectivity classes. ∎
@@
Since $\mathcal{V}$ is a $\lambda$-quasivariety we can write the unit $I$ as a regular quotient of a coproduct of $\lambda$-presentable regular projective objects. Moreover, since $I$ is $\lambda$-presentable, we can assume such coproduct to be $\lambda$-small. But $\lambda$-small coproducts of $\lambda$-presentable regular projective objects is still $\lambda$-presentable regular projective. It follows that there exists a $\lambda$-presentable and regular projective object $I^{*}$ together with a regular epimorphism
$e\colon I^{*}\twoheadrightarrow I.$
We shall now fix such $I^{*}$ and $e$ and, without loss of generality, assume that $I^{*}\in\mathcal{P}$ (one can always enlarge $\mathcal{P}$ to contain $I^{*}$).
@@
Given $I^{*}$ as above, by pre-composing with $e$, every morphism in a $\mathcal{V}$-category $\mathcal{K}$ can be seen as a $I^{*}$-morphism.
As it happens in the case of ordinary morphisms (Definition 8.6), $I^{*}$-morphisms can be composed with $P$-morphisms for any $P\in\mathcal{P}$. Indeed, given $P\in\mathcal{P}$, tensoring with $e$ gives a regular epimorphism
$e_{P}\colon P\otimes I^{*}\to P$
which, since $P$ is regular projective, splits giving $s_{e}\colon P\to P\otimes I^{*}$ such that $e_{P}s_{P}=1_{P}$. Now, given a $I^{*}$-morphism $(g,I^{*})\colon X\to Y$ and a $P$-morphism $(f,P)\colon Y\to X$, we define their composite as the $P$-morphism
$(fg,P)\colon P\stackrel{{\scriptstyle s_{e}}}{{\longrightarrow}}P\otimes I^{*}\stackrel{{\scriptstyle f\otimes g}}{{\xrightarrow{\hskip 22.76228pt}}}\mathcal{K}(Y,Z)\otimes\mathcal{K}(X,Y)\stackrel{{\scriptstyle\circ}}{{\longrightarrow}}\mathcal{K}(X,Z).$
Finally, note that the composite of a a morphism $g\colon X\to Y$ with a $P$-morphism $(f,P)\colon Y\to Z$ with is the same as the composite of the $I^{*}$-morphism $(ge,I^{*})$, induced by $g$ through $e$, with $(f,P)$; this is because $s_{P}$ is a section of $e_{P}$.
@@
A morphism $s\colon X\to Y$ in a $\mathcal{V}$-category $\mathcal{K}$ is an $\mathcal{E}$-split monomorphism if and only if there exists $(t,I^{*})\colon Y\to X$ for which
$(ts,I^{*})=(1_{X},I^{*})$
in the sense of Definition 8.6.
@@
is surjective. By taking an element that is mapped to $(1_{X},I^{*})$ we find $(t,I^{*})\colon Y\to X$ as in the statement.
Conversely, suppose that there exist $(t,I^{*})\colon Y\to X$ for which $(ts,I^{*})=(1_{X},I^{*})$; in the setting of Remark 8.18 the equality can be rewritten as
$(t(se),I^{*})=(1_{X},I^{*})$
where $(se,I^{*})$ is the $I^{*}$-morphism induced by precomposing $s\colon I\to\mathcal{K}(X,Y)$ with $e\colon I^{*}\to I$. To prove that $s$ is an $\mathcal{E}$-split monomorphism it is enough to show that
$\mathcal{V}_{0}(P,\mathcal{K}(s,X))\colon\mathcal{V}_{0}(P,\mathcal{K}(Y,X))\to\mathcal{V}_{0}(P,\mathcal{K}(X,X))$
is surjective for any $P\in\mathcal{P}$. Given $(f,P)\colon X\to X$ in the codomain, we can consider the $P$-morphism $(ft,P)\colon Y\to X$ defined in Remark 8.18 above. By definition, $\mathcal{V}_{0}(P,\mathcal{K}(s,X))$ sends $(ft,P)$ to the the composite
$(ft,P)\circ(s,I)$ $=(ft,P)\circ(se,I^{*})$
$=(f,P)\circ(t(se),I^{*})$
$=(f,P)\circ(1_{X},I^{*})$
$=(f,P),$
where the first composition is in the sense of Definition 8.6 and the others are in the sense of Remark 8.18. It follows that $\mathcal{V}_{0}(P,\mathcal{K}(s,X))$is surjective, and thus $\mathcal{K}(s,X)$ is a regular epimorphism and $s$ an $\mathcal{E}$-split monomorphism. ∎
@@
(1) follows from Corollary 5.4 and (2) is a consequence of Proposition 4.9 since monomorphisms are closed under $\lambda$-filtered colimits in $\mathcal{V}^{\to}$.
(3) We apply the same proof of [2, Proposition 2.30] with some changes where the notion of $\mathcal{E}$-split morphism is needed.
Given a $(\lambda,\mathcal{E})$-pure morphism $f\colon A\to B$, we can write it as a $\lambda$-filtered colimit in $\mathcal{K}^{\to}$ of maps $f_{i}\colon A_{i}\to B_{i}$ between $\lambda$-presentable objects with connecting morphisms $(u_{i},v_{i})\colon f_{i}\to f$. As usual, all the morphisms $u_{i}$ and $v_{i}$ can be seen as $I^{*}$-morphisms (by pre-composing with $e\colon I^{*}\to I$), giving commutative squares as below.
@@
It is shown in [2, Proposition 2.30] that the colimit of the $\bar{f}_{i}$ is still $f$; thus we only need to show that each $\bar{f}_{i}$ is $\mathcal{E}$-split.
We can see the identity on $A$ as an $I^{*}$-morphism and hence, acting by transposition, as a map $\bar{1}_{A}\colon A\to I^{*}\pitchfork A$; similarly each $(t_{i},I^{*})$ corresponds to a morphism $\bar{t}_{i}\colon B_{i}\to I^{*}\pitchfork A$. The fact that $(t_{i}f,I^{*})=(u_{i},I^{*})$ means that the square below commutes.